# subfigure0

### v1

@@
We study the effect of supersymmetry breaking on Krylov complexity in the $\mathcal{N}=2$ SYK model under irrelevant and mass deformations of the Hamiltonian. The irrelevant deformation breaks $\mathcal{N}=2$ supersymmetry down to $\mathcal{N}=1$, while the mass deformation breaks supersymmetry completely. Using Krylov subspace methods, we analyze the Lanczos sequence, Krylov dimension, complexity, and entropy at finite system size as functions of deformation strength. Both deformations enlarge the Krylov space, but the mass deformation has a stronger effect. Krylov complexity exhibits initial quadratic growth, followed by linear growth across both deformations. We observe that both deformations increase the quadratic and linear growth rates of Krylov complexity at early times. At late times, the irrelevant deformation increases the saturation complexity as a fraction of the Krylov dimension, while the mass deformation decreases it. This reveals distinct signatures of how supersymmetry breaking impacts quantum complexity.
@@ 1 Introduction
Krylov complexity has recently emerged as a powerful tool for studying quantum chaos in many-body quantum systems and gravity [Nandy_2024, Rabinovici_2025, Baiguera_2025]. Roughly, Krylov complexity measures how quickly an operator grows under generic Hamiltonian evolution. In a chaotic system, Krylov complexity typically grows exponentially at early times, which provides a useful probe of quantum chaos. The notion of Krylov complexity grew out of a need to understand quantum chaos beyond the semiclassical limit. Semiclassically, chaos is often diagnosed through the Lyapunov exponent, which quantifies the exponential growth of out-of-time-order correlators. The Lyapunov exponent is bounded by the temperature of the system, and maximally chaotic systems, such as black holes, saturate this bound [MSS_2016, Tsuji_2018]. Krylov complexity extends this idea to more general settings, quantifying operator growth under generic, non-integrable Hamiltonian dynamics. When the Lyapunov exponent can be defined, it is bounded by the exponential growth rate of Krylov complexity [Parker_2019].
@@
The fundamental idea behind Krylov complexity, and Krylov subspace methods generally, is to identify a minimal subspace in which an operator evolves [Liesen_2012]. This subspace is constructed by means of the Lanczos algorithm [Cullum_2002]. In this algorithm, an initial operator is evolved using the Liouvillian, the generator of time evolution. After each step, the overlap with all previous operators is subtracted in a Gram-Schmidt-like procedure. The resulting orthonormal basis is called the Krylov basis. In addition to the Krylov basis, the Lanczos algorithm generates a sequence of coefficients, called the Lanczos coefficients, which correspond to the size of the Krylov basis elements. In a chaotic system, the Lanczos coefficients initially grow in a system-dependent way before entering a regime of linear growth at intermediate times. The coefficients then plateau and finally decay [Parker_2019]. The Lanczos sequence captures the complete system dynamics, from which Krylov complexity can be computed.
@@
One interesting model which has been used to study Krylov complexity is the Sachdev-Ye-Kitaev (SYK) model, a quantum mechanical model of interacting fermions with random couplings [Sachdev_1993, kitaev2015talks, Maldacena_2016, Polchinski_2016, Garc_a_Garc_a_2016, Rosenhaus_2019, Berkooz_2024]. Because its Lyapunov exponent saturates the bound, the SYK model is maximally chaotic and has been proposed as a simple, one-dimensional model of quantum black holes. At low energies, SYK exhibits emergent conformal symmetry. In the soft sector, the model can be described holographically by Jackiw-Teitelboim (JT) gravity, a two-dimensional gravitational theory in anti-de Sitter space [Jensen_2016, Mertens_2023, Turiaci_2024, Saad_2019]. The Lanczos sequence of the model was shown to exhibit initial linear growth characteristic of chaotic systems [Parker_2019]. Krylov complexity also grows exponentially at early times. When the number of interactions is large, the exponential growth rate of Krylov complexity matches the Lyapunov exponent, again suggesting maximal chaos.
@@
One major way to generalize the SYK model is to introduce supersymmetry, which can be used to study supersymmetric black holes [Fu_2017, Berkooz_2020]. There are different versions of supersymmetric SYK depending on the number of supercharges. Roughly, supersymmetric SYK models possess the same form as their non-supersymmetric counterparts, but the couplings are no longer independent. These correlations change the structure of the large-$N$ equations and the conformal dimension of the fermions, but generally the models are maximally chaotic in the sense that their Lyapunov exponents saturate the bound [Peng_2017]. Starting at $\mathcal{N}=2$, which possesses two complex supercharges, the model develops a large degeneracy of ground states with exactly zero energy, called BPS states [Heydeman_2024, Heydeman_2023]. Between the discrete BPS states and non-BPS continuum, there is a gap of order $1/S$, where $S$ is the entropy. Because of its interesting properties, $\mathcal{N}=2$ SYK has been used to study the thermodynamics of extremal, supersymmetric black holes in higher dimensions.
@@
Although the Lyapunov exponent of supersymmetric SYK in the large-$N$ limit has been studied extensively, Krylov complexity of supersymmetric SYK has received less attention. A study of Krylov complexity of $\mathcal{N}=1$ SYK has been conducted, but it was embedded within a larger study of $T\bar{T}$-deformed SYK that also analyzed other measures of chaos [He_2022]. The chaotic properties of $\mathcal{N}=2$ have also been studied, but primarily within the context of random matrix theory [Kanazawa_2017]. The related concept of spread complexity has also been applied to supersymmetric systems, but not to $\mathcal{N}=2$ SYK [Das2025]. This paper is devoted to a thorough study of Krylov complexity of the $\mathcal{N}=2$ SYK model. Various deformations are explored that break the supersymmetry, which enable the effects of supersymmetry on Krylov complexity to be studied.
@@
The first deformation that is treated is a UV deformation introduced in [Heydeman_2024]. Because this deformation is irrelevant, it does not spoil the IR physics, which corresponds to the gravitational dynamics. The deformation breaks $\mathcal{N}=2$ down to $\mathcal{N}=1$, eliminating the BPS states and lifting the ground-state degeneracy for even parametrically small values of the deformation parameter. This deformation can be seen as interpolating between two types of black holes [Iliesiu_2021, Heydeman_2021]. The first corresponds to extremal, supersymmetric black holes which carry a large entropy reflecting the ground-state degeneracy. The second type of black hole corresponds to near-extremal, non-supersymmetric black holes which display no ground-state degeneracy and an average energy spacing that is exponentially suppressed in the entropy.
@@
While the UV deformation is interesting on the level of black hole physics, this paper studies the effect of the deformation on Krylov complexity at finite $N$ and infinite temperature. Some relevant questions include the following. How does supersymmetry breaking alter the Lanczos sequence? How do the coefficients grow initially? How do the sequences compare across different operators and system sizes? How does Krylov complexity respond to the deformation? Does it grow exponentially at early times, or does it follow another growth profile? These questions will be answered numerically, which presents significant computational challenges. These challenges include the exponential growth of the Hilbert space with system size and the breakdown of orthogonality in the Krylov basis as the number of iterations increases. Addressing these challenges necessitates the use of optimized, parallelized implementations of the Lanczos algorithm.
@@
The second deformation that is explored is a relevant deformation that corresponds on the gravity side to deformations away from two-dimensional anti-de Sitter space. There are many types of relevant deformations that can be considered, but one of the simplest involves adding the Hamiltonian of two-body SYK with complex fermions, which is a free, integrable theory [Anninos_2023, Chapman_2024]. This mass deformation provides another pattern of supersymmetry breaking that breaks $\mathcal{N}=2$ directly down to $\mathcal{N}=0$. The BPS degeneracy is lifted, and the energies are no longer bounded from below by zero. The incorporation of an integrable Hamiltonian helps steer the model toward integrability, which has a direct bearing on Krylov complexity that distinguishes between chaotic and integrable dynamics [Garcia_Garcia_2018, Rabinovici_2022, Rabinovici_2022_2]. The same types of questions as before will be addressed.
@@ 2.1 $\mathcal{N}=0$ SYK
where $\{c_{i},c_{j}^{\dagger}\}=\delta_{ij}$ and $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$. The total fermion number can be defined as $F\equiv\sum_{i=1}^{N/2}c_{i}^{\dagger}c_{i}$. This is not a conserved charge, but fermion number parity, $F\text{ mod }2$, is conserved. This entails that the Hamiltonian can be block-diagonalized into sectors of even and odd fermion number [Cotler_2017]. In practice, one way to implement Dirac fermions is by means of the Jordan-Wigner transformation, which maps them to a chain of spin-1/2 particles. The construction is
@@
The $\mathcal{N}=0$ SYK model involving Dirac fermions, called complex SYK, is built in much the same way as the Majorana version [Gu_2020, Pethybridge_2024, Zhang_2025, Sachdev_2015]. This model can be defined for any $q$-body interaction, but for $q=4$ the Hamiltonian takes the form
@@
One way of investigating the chaotic properties of the SYK model is to study its energy eigenvalues, or spectral properties. Figure 1(a) shows the spectral density $\rho(E)$ of $\mathcal{N}=0$ SYK with $q=4$ interactions. The distribution was obtained by exactly diagonalizing 500 samples of the Hamiltonian. The number of fermions is $N=10$, and the model was diagonalized in the sector $F=5$ with $J=1$. The density resembles the Wigner semicircle distribution, which arises from the eigenvalues of random matrices [Kravtsov_2012, Livan_2018]. The distribution does not peak sharply at $E=0$ and drops off rapidly at large energies. The shape of the distribution results from repulsion between energy levels, which is characteristic of chaotic systems.
A further way of investigating the spectral properties of the model is to examine the density of spacings between successive eigenvalues $p(s)$, called the level spacings. This is shown in Figure 2(a). Eigenvalue repulsion causes the distribution to peak at $s>0$. Besides the peak, the gap ratio $r$ is another way of characterizing the level spacing distribution, given by
@@
$H=\sum_{ij}J_{ij}\,c_{i}^{\dagger}c_{j}$ (2.8)
where $J_{ij}$ is a Hermitian matrix whose elements are sampled from a Gaussian distribution with zero mean and variance $J^{2}/4N$. This model is a free, integrable theory, which can again be seen by studying its spectral properties. Figure 1(b) shows the spectral density for the $q=2$ model. The distribution is shaped like a Gaussian, peaking sharply at $E=0$ and tapering off smoothly at large energies. Figure 2(b) shows the distribution of level spacings. Because the eigenvalues do not repel each other, the level spacings follow a Poisson distribution that peaks at $s=0$, where eigenvalues cluster together. The mean gap ratio is $\langle r\rangle\approx 0.38$, which is characteristic of integrable systems.
@@ 2.2 $\mathcal{N}=1,2$ SYK
The SYK model can be generalized to include supersymmetry, which can be used to study supersymmetric black holes [Fu_2017, Peng_2017, Kanazawa_2017, Peng_2021]. There are various types of supersymmetric SYK models corresponding to different numbers of supercharges, but two of the most studied are $\mathcal{N}=1$ and $\mathcal{N}=2$. The $\mathcal{N}=1$ model uses Majorana fermions, while $\mathcal{N}=2$ uses Dirac fermions. In $\mathcal{N}=1$, the supercharge $Q$ is defined as
@@
At finite $N$, the $\mathcal{N}=1$ model has a non-zero ground state energy, which breaks supersymmetry, but it is recovered in the large-$N$ limit [Fu_2017]. Notice that the Hamiltonian of Equation (2.10) has the same form as the Hamiltonian of $\mathcal{N}=0$ Majorana SYK with $\tilde{q}$ interactions, but the couplings $J_{i_{1}\cdots i_{\tilde{q}}}$ are not independent. Instead, they roughly correspond to the covariance matrix of the couplings $C_{i_{1}\cdots i_{q}}$. For $q=3$, these couplings are given by
@@
Like the complex SYK model, the $\mathcal{N}=2$ model possesses a global $U(1)$ symmetry, which leads to conservation of total fermion number $F$. The $\mathcal{N}=2$ theory possesses an additional $U(1)$ R-symmetry generated by the operator
$R=\frac{1}{2q}\sum_{i}[c_{i}^{\dagger},c_{i}]$ (2.14)
@@
While the $U(1)_{R}$ charge does not commute with the supercharges, there is a $Z_{q}$ subgroup of this symmetry that does, which transforms the fermions as $c_{i}\to e^{\frac{2\pi in}{q}}c_{i}$, $n\in\mathbb{Z}_{q}$ [Fu_2017].
Because the $U(1)_{R}$ charge commutes with the Hamiltonian, the Hilbert space can be decomposed as $\mathcal{H}=\bigoplus_{k}\mathcal{H}_{k}$, where $\mathcal{H}_{k}$ is the subspace of states with $U(1)_{R}$ charge $k$. In each space $\mathcal{H}_{k}$, there are BPS and non-BPS supermultiplets. BPS states are zero-energy states annihilated by the supercharges $Q$ and $Q^{\dagger}$ [Chen2024, Chang2024]. Non-BPS states are raised or lowered to other $R$-charge sectors by the supercharges. BPS states appear with a large degeneracy at $E=0$. Non-BPS states exhibit a dense spectrum starting from a non-zero energy.

### v2

@@
We study the effect of supersymmetry breaking on Krylov complexity in the $\mathcal{N}=2$ SYK model under irrelevant and mass deformations of the Hamiltonian. The irrelevant deformation breaks $\mathcal{N}=2$ supersymmetry down to $\mathcal{N}=1$, while the mass deformation breaks supersymmetry completely. Using Krylov subspace methods, we analyze the Lanczos sequence, Krylov dimension, complexity, and entropy at finite system size as functions of deformation strength. Both deformations enlarge the Krylov space, but the mass deformation, which completely lifts the energy degeneracy, generates a stronger enhancement. Krylov complexity exhibits initial quadratic growth, followed by linear growth, across both deformations. We observe that both deformations increase the quadratic and linear growth rates of Krylov complexity at early times. At late times, the irrelevant deformation increases the saturation complexity as a fraction of the Krylov dimension, while the mass deformation decreases it. This reveals distinct signatures of how supersymmetry breaking impacts quantum complexity.
@@ 1 Introduction
Krylov complexity has recently emerged as a powerful tool for studying quantum chaos in many-body quantum systems and gravity [50, 60, 10]. Roughly, Krylov complexity measures how quickly an operator grows under generic Hamiltonian evolution. In a chaotic system, Krylov complexity typically grows exponentially at early times, which provides a useful probe of quantum chaos. The notion of Krylov complexity grew out of a need to understand quantum chaos beyond the semiclassical limit. Semiclassically, chaos is often diagnosed through the Lyapunov exponent, which quantifies the exponential growth of out-of-time-order correlators. The Lyapunov exponent is bounded by the temperature of the system, and maximally chaotic systems, such as black holes, saturate this bound [47, 66]. Krylov complexity extends this idea to more general settings, quantifying operator growth under generic, non-integrable Hamiltonian dynamics. When the Lyapunov exponent can be defined, it is bounded by the exponential growth rate of Krylov complexity [51].
@@
The fundamental idea behind Krylov complexity, and Krylov subspace methods generally, is to identify a minimal subspace in which an operator evolves [45]. This subspace is constructed by means of the Lanczos algorithm [24]. In this algorithm, an initial operator is evolved using the Liouvillian, the generator of time evolution. After each step, the overlap with all previous operators is subtracted in a Gram-Schmidt-like procedure. The resulting orthonormal basis is called the Krylov basis. In addition to the Krylov basis, the Lanczos algorithm generates a sequence of coefficients, called the Lanczos coefficients, which correspond to the size of the Krylov basis elements. In a chaotic system, the Lanczos coefficients initially grow in a system-dependent way before entering a regime of linear growth at intermediate times. The coefficients then plateau and finally decay [51]. The Lanczos sequence captures the complete system dynamics, from which Krylov complexity can be computed.
@@
One interesting model which has been used to study Krylov complexity is the Sachdev-Ye-Kitaev (SYK) model, a quantum mechanical model of interacting fermions with random couplings [64, 43, 48, 55, 28, 62, 14]. Because its Lyapunov exponent saturates the bound, the SYK model is maximally chaotic and has been proposed as a simple, one-dimensional model of quantum black holes. At low energies, SYK exhibits emergent conformal symmetry. In the soft sector, the model can be described holographically by Jackiw-Teitelboim (JT) gravity, a two-dimensional gravitational theory in anti-de Sitter space [39, 49, 67, 63]. The Lanczos sequence of the model was shown to exhibit initial linear growth characteristic of chaotic systems [51]. Krylov complexity also grows exponentially at early times. When the number of interactions is large, the exponential growth rate of Krylov complexity matches the Lyapunov exponent, again suggesting maximal chaos.
@@
One major way to generalize the SYK model is to introduce supersymmetry, which can be used to study supersymmetric black holes [26, 13]. There are different versions of supersymmetric SYK depending on the number of supercharges. Roughly, supersymmetric SYK models possess the same form as their non-supersymmetric counterparts, but the couplings are no longer independent. These correlations change the structure of the large-$N$ equations and the conformal dimension of the fermions, but generally the models are maximally chaotic in the sense that their Lyapunov exponents saturate the bound [52]. Starting at $\mathcal{N}=2$, which possesses two complex supercharges, the model develops a large degeneracy of ground states with exactly zero energy, called BPS states [37, 35]. Between the discrete BPS states and non-BPS continuum, there is a gap of order $1/S$, where $S$ is the entropy. Because of its interesting properties, $\mathcal{N}=2$ SYK has been used to study the thermodynamics of extremal, supersymmetric black holes in higher dimensions.
@@
Although the Lyapunov exponent of supersymmetric SYK in the large-$N$ limit has been studied extensively, Krylov complexity of supersymmetric SYK has received less attention. A study of Krylov complexity of $\mathcal{N}=1$ SYK has been conducted, but it was embedded within a larger study of $T\bar{T}$-deformed SYK that also analyzed other measures of chaos [32]. The chaotic properties of $\mathcal{N}=2$ have also been studied, but primarily within the context of random matrix theory [41]. The related concept of spread complexity has also been applied to supersymmetric systems, but not to $\mathcal{N}=2$ SYK [25]. This paper is devoted to a thorough study of Krylov complexity of the $\mathcal{N}=2$ SYK model. Various deformations are explored that break the supersymmetry, which enable the effects of supersymmetry on Krylov complexity to be studied.
@@
The first deformation that is treated is a UV deformation introduced in [37]. Because this deformation is irrelevant, it does not spoil the IR physics, which corresponds to the gravitational dynamics. The deformation breaks $\mathcal{N}=2$ down to $\mathcal{N}=1$, eliminating the BPS states and lifting the ground-state degeneracy for even parametrically small values of the deformation parameter. This deformation can be seen as interpolating between two types of black holes [38, 36]. The first corresponds to extremal, supersymmetric black holes which carry a large entropy reflecting the ground-state degeneracy. The second type of black hole corresponds to near-extremal, non-supersymmetric black holes which display no ground-state degeneracy and an average energy spacing that is exponentially suppressed in the entropy.
@@
While the UV deformation is interesting on the level of black hole physics, this paper studies the effect of the deformation on Krylov complexity at finite $N$ and infinite temperature. The way that supersymmetry breaking modifies the Lanczos sequence is examined across a range of different operators and system sizes. The response of Krylov complexity to the deformation is analyzed, including its early-time behavior and potential departures from exponential growth. The late-time saturation of Krylov complexity is also studied. The analysis is performed numerically, which presents significant computational challenges. These challenges include the exponential growth of the Hilbert space with system size and the breakdown of orthogonality in the Krylov basis as the number of iterations increases. Addressing these challenges necessitates the use of optimized, parallelized implementations of the Lanczos algorithm.
@@
The second deformation that is explored is a relevant deformation that corresponds on the gravity side to deformations away from two-dimensional anti-de Sitter space. There are many types of relevant deformations that can be considered, but one of the simplest involves adding the Hamiltonian of two-body SYK with complex fermions, which is a free, integrable theory [8, 20]. This mass deformation provides another pattern of supersymmetry breaking that breaks $\mathcal{N}=2$ directly down to $\mathcal{N}=0$. The BPS degeneracy is lifted, and the energies are no longer bounded from below by zero. The incorporation of an integrable Hamiltonian helps steer the model toward integrability, which has a direct bearing on Krylov complexity that distinguishes between chaotic and integrable dynamics [27, 57, 58]. The same kind of analysis as before is performed.
@@ 2.1 $\mathcal{N}=0$ SYK
where $\{c_{i},c_{j}^{\dagger}\}=\delta_{ij}$ and $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$. The total fermion number can be defined as $F\equiv\sum_{i=1}^{N/2}c_{i}^{\dagger}c_{i}$. This is not a conserved charge, but fermion number parity, $F\text{ mod }2$, is conserved. This entails that the Hamiltonian can be block-diagonalized into sectors of even and odd fermion number [22]. In practice, one way to implement Dirac fermions is by means of the Jordan-Wigner transformation, which maps them to a chain of spin-1/2 particles. The construction is
@@
The $\mathcal{N}=0$ SYK model involving Dirac fermions, called complex SYK, is built in much the same way as the Majorana version [29, 54, 69, 65]. This model can be defined for any $q$-body interaction, but for $q=4$ the Hamiltonian takes the form
@@
One way of investigating the chaotic properties of the SYK model is to study its energy eigenvalues, or spectral properties. Figure 1(a) shows the spectral density $\rho(E)$ of $\mathcal{N}=0$ SYK with $q=4$ interactions. The distribution was obtained by exactly diagonalizing 500 samples of the Hamiltonian. The number of fermions is $N=10$, and the model was diagonalized in the sector $F=5$ with $J=1$. The density resembles the Wigner semicircle distribution, which arises from the eigenvalues of random matrices [44, 46]. The distribution does not peak sharply at $E=0$ and drops off rapidly at large energies. The shape of the distribution results from repulsion between energy levels, which is characteristic of chaotic systems.
A further way of investigating the spectral properties of the model is to examine the density of spacings between successive eigenvalues $p(s)$, called the level spacings. This is shown in Figure 2(a). The density is a Wigner-Dyson distribution, which peaks at $s>0$ due to eigenvalue repulsion. Besides the peak, the gap ratio $r$ is another way of characterizing the level spacing distribution, given by
@@
$H=\frac{1}{2}\sum_{ij}J_{ij}\,c_{i}^{\dagger}c_{j}$ (2.8)
where $J_{ij}$ is a Hermitian matrix whose elements are sampled from a Gaussian distribution with zero mean and variance $J^{2}/N$. This model is a free, integrable theory, which can again be seen by studying its spectral properties. Figure 1(b) shows the spectral density for the $q=2$ model. The distribution is shaped like a Gaussian, peaking sharply at $E=0$ and tapering off smoothly at large energies. Figure 2(b) shows the distribution of level spacings. Because the eigenvalues do not repel each other, the level spacings follow a Poisson distribution that peaks at $s=0$, where eigenvalues cluster together. The mean gap ratio is $\langle r\rangle\approx 0.38$, which is characteristic of integrable systems.
@@ 2.2 $\mathcal{N}=1,2$ SYK
The SYK model can be generalized to include supersymmetry, which can be used to study supersymmetric black holes [26, 52, 41, 53]. There are various types of supersymmetric SYK models corresponding to different numbers of supercharges, but two of the most studied are $\mathcal{N}=1$ and $\mathcal{N}=2$. The $\mathcal{N}=1$ model uses Majorana fermions, while $\mathcal{N}=2$ uses Dirac fermions. In $\mathcal{N}=1$, the supercharge $Q$ is defined as
@@
At finite $N$, the $\mathcal{N}=1$ model has a non-zero ground state energy, which breaks supersymmetry, but it is recovered in the large-$N$ limit [26]. Notice that the Hamiltonian of Equation (2.10) has the same form as the Hamiltonian of $\mathcal{N}=0$ Majorana SYK with $\tilde{q}$ interactions, but the couplings $J_{i_{1}\cdots i_{\tilde{q}}}$ are not independent. Instead, they roughly correspond to the covariance matrix of the couplings $C_{i_{1}\cdots i_{q}}$. For $q=3$, these couplings are given by
@@
Like the complex SYK model, the $\mathcal{N}=2$ model possesses a global $U(1)$ symmetry, which leads to conservation of total fermion number $F$. The $\mathcal{N}=2$ theory possesses a related $U(1)$ R-symmetry generated by the operator
$R=\frac{1}{2q}\sum_{i}[c_{i}^{\dagger},c_{i}]=\frac{F}{q}-\frac{N}{2q}$ (2.14)
@@
While the $U(1)_{R}$ charge does not commute with the supercharges, there is a $Z_{q}$ subgroup of this symmetry that does, which transforms the fermions as $c_{i}\to e^{\frac{2\pi in}{q}}c_{i}$, $n\in\mathbb{Z}_{q}$ [26].
Because the $U(1)_{R}$ charge commutes with the Hamiltonian, the Hilbert space can be decomposed as $\mathcal{H}=\bigoplus_{k}\mathcal{H}_{k}$, where $\mathcal{H}_{k}$ is the subspace of states with $U(1)_{R}$ charge $k$. In each space $\mathcal{H}_{k}$, there are BPS and non-BPS supermultiplets. BPS states are zero-energy states annihilated by the supercharges $Q$ and $Q^{\dagger}$ [21, 19]. Non-BPS states are raised or lowered to other $R$-charge sectors by the supercharges. BPS states appear with a large degeneracy at $E=0$. Non-BPS states exhibit a dense spectrum starting from a non-zero energy.